It says to solve the system of equations using the substitution method.

I think I'm supposed to see what X and Y are equal to, but I keep getting some pretty messed up answers.

Answer by LilSkittleMd(119)

x+y=1
3x+4y=1
Solve for x or y
x=1-y
Substitute this into the second equation
3(1-y)+4y=1
3-3y+4y=1
3+y=1
y=-2
x=1+2
x=3
(3,-2)
x=3
y=-2